OpenStudy (maddss):

The center and vertices of the hyperbola y^2-4x^2-6y-8x-11=0 are: A) center (3, -1) vertices (-1, 1) and (7, -1) B) center (-1, 3) vertices (-1, 7) and (-1, -1) C) center (3, 1) vertices (5, 1) and (1, 1) D) center (3, 1) vertices (3, 5) and (3, 3)

OpenStudy (tkhunny):

Complete the square in both x and y.

OpenStudy (maddss):

what?

OpenStudy (tkhunny):

You're saying you've never "Completed the Square"? y^2-4x^2-6y-8x-11=0 Rearrange y^2 - 6y - 4x^2 - 8x - 11 = 0 Regroup and set up (y^2 - 6y + ____) - 4(x^2 + 2x + ____) = 11 + ____ - 4(____) Fill in the blanks!

Directrix (directrix):

@maddss Did you see the graph for this? Maybe @tkhunny will give a solution. http://openstudy.com/users/maddss#/updates/514e1555e4b0ae0b658aed6b